OpenStudy (anonymous):

Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4x-2.

OpenStudy (anonymous):

Do you need help ?

OpenStudy (anonymous):

s(x) + s(x) * a(x) = Total production ^--- starting with this no matter what increasing by a(x) factor on what he can produce at first so s(x) * a(x)

OpenStudy (anonymous):

Let P(x) be the production function per square root for x>=1 P(x)=s(x)+s(x)*a(x) = 10 + 10*(1.4x-2)

OpenStudy (anonymous):

x is the number of square feet you want to maintain

OpenStudy (anonymous):

so when it asks how much he will maintain for 5, that means he's maintaining 4 more square feet

OpenStudy (anonymous):

10+10(1.4(4)-2)=Production